# thanks: These authors contributed equallythanks: These authors contributed equally

### v1

@@ Abstract
Discussion.— We have proposed a new type of valley dependent topological phenomena, dubbed the ZBC-QVHE, induced by the VEN. Although the Euler number is a topological invariant of two bands, the helical DW states are stable even in the presence of additional bulk bands as long as they do not disturb the half-integral Euler number of Euler bands at each valley. This is because the crossing between helical DW states from each valley is further protected by $M_{y}$ symmetry of the DWC, which is well-defined even in multi-band systems. In the case of the helical DW states protected by ~$\tilde{\mathcal{I}}^{\prime}_{\text{ST}}$ and $\tilde{\Pi}$ symmetries, the stability of the helical DW state crossing is guaranteed by the fact that $H_{\text{DW}}$ has non-zero $\mathcal{Z}_{2}$ 00-dimensional charge when commuting ~$\tilde{\mathcal{I}}^{\prime}_{\text{ST}}$ and $\tilde{\Pi}$ symmetries are present [67].
In addition to h-BX materials, we have identified various candidate systems where ZBC-QVHE can be realized such as large-angle twisted bilayer graphene [29, 70, 71, 72, 73, 74, 75], and hexagonal bilayer spin systems [76, 77, 78, 79, 80, 81] as described in detail in SM.
Acknowledgements R.G. and C.M. thank Yuting Qian for fruitful discussions. R.G, C.M, S.K, and B-J.Y were supported by Samsung Science and Technology Foundation under Project Number SSTF-BA2002-06, the National Research Foundation of Korea (NRF) grant funded by the Korean government (MSIT) (No.2021R1A2C4002773, and No. NRF-2021R1A5A1032996).

### v2

@@ Abstract
Discussion.— We have proposed a new type of valley dependent topological phenomena, dubbed the ZBC-QVHE, induced by the VEN. Although the Euler number is a topological invariant of two bands, the helical DW states are stable even in the presence of additional bulk bands as long as they do not disturb the half-integral Euler number of Euler bands at each valley. This is because the crossing between helical DW states from each valley is further protected by $M_{y}$ symmetry of the DWC, which is well-defined even in multi-band systems. In the case of the helical DW states protected by ~$\tilde{\mathcal{I}}^{\prime}_{\text{ST}}$ and $\tilde{\Pi}$ symmetries, the stability of the helical DW state crossing is guaranteed by the fact that $H_{\text{DW}}$ has non-zero $\mathcal{Z}_{2}$ 00-dimensional charge when commuting ~$\tilde{\mathcal{I}}^{\prime}_{\text{ST}}$ and $\tilde{\Pi}$ symmetries are present Bzdušek and Sigrist (2017).
In addition to h-BX materials, we have identified various candidate systems where ZBC-QVHE can be realized such as large-angle twisted bilayer graphene Mondal et al. (2023); Mele (2012); Shallcross et al. (2008); Mele (2010); Kindermann (2015); Park et al. (2019); Liu et al. (2022), and hexagonal bilayer spin systems Wang et al. (2019); Islam and Benjamin (2016); Yang et al. (2021); Ezawa (2015); Li et al. (2013); Rakhmanov et al. (2012) as described in detail in SM. Furthermore, we anticipate that ZBC QVHE and its DW modes can be readily achieved in classical wave systems, where many QVHE phenomena have already been observed  Lu et al. (2017); Li et al. (2022); Lu et al. (2018, 2018); Yang et al. (2021).
Finally, let us discuss the stability of the DW modes under various perturbations. First, we examined the DW $M_{y}$ symmetry-breaking effect by considering the unequal lengths of two domains as well as the vertical and lateral shifts of domains. By performing first-principle calculations, we confirmed the stability of the DW modes against $M_{y}$ symmetry breaking when the width of each domain is large enough and domain mismatch is within the experimentally accessible window Note (1). Furthermore, we considered local disorder effects induced by random onsite potentials and local lattice distortions near the DW by analyzing finite-size lattices and performing the spectral function analysis Schubert et al. (2012); Wu et al. (2013); Queiroz and Schnyder (2015); Pieper and Fehske (2016); Araújo and Andrade (2019); Ghadimi et al. (2021). As expected, in the presence of disorder, since the mirror symmetry is broken, the DW states lose their topological protection. However, we found that even when the disorder potential strength is as large as the band gap, the only influence of local disorder effect is at most to open tiny gaps at the crossing of helical DW states. Thus, our numerical study demonstrates the stability of the helical DW states against local perturbations, and their experimental feasibility Note (1).
Acknowledgements We thank Yuting Qian for the fruitful discussions. R.G, C.M, S.K, and B-J.Y were supported by Samsung Science and Technology Foundation under Project Number SSTF-BA2002-06 and the National Research Foundation of Korea (NRF) grant funded by the Korean government (MSIT) No. NRF-2021R1A5A1032996. This research was supported by GRDC (Global Research Development Center) Cooperative Hub Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Science and ICT(MSIT) (RS-2023-00258359).